# A spectral method for multi-view subspace learning using the product of projections

### v1

@@ A.1. Proofs of main results
$\leq\frac{\dim(\mathcal{J})^{1/2}\left\|P_{\operatorname*{col}(X_{1})}\Delta_{2}+\Delta_{1}P_{\operatorname*{col}(X_{2})}+\Delta_{1}\Delta_{2}+\Delta_{2}P_{\operatorname*{col}(X_{1})}+P_{\operatorname*{col}(X_{2})}\Delta_{1}+\Delta_{2}\Delta_{1}\right\|_{2}}{1-\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})}$
@@
$=\frac{\dim(\mathcal{J})^{1/2}\left\|R_{\mathcal{J}}+R_{\mathcal{J}}^{\top}\right\|_{2}}{1-\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})}.$
Individual components. Since $\mathcal{J}\perp\mathcal{I}_{1}$ and $\mathcal{J}\perp\mathcal{I}_{2}$, we have that $P_{\mathcal{J}^{\perp}}P_{\operatorname*{col}(X_{1})}=P_{\mathcal{I}_{1}}$ and $P_{\mathcal{J}^{\perp}}P_{\operatorname*{col}(X_{2})}=P_{\mathcal{I}_{2}}$. Therefore, we have that $P_{\mathcal{J}^{\perp}}P_{\operatorname*{col}(X_{k})}$ has exactly $r_{\mathcal{I}_{k}}$ eigenvalues equal to $1$ and the rest equal to 00. Therefore, applying Davis-Kahan, we get:
$\|P_{\widehat{\mathcal{I}}_{k}}-P_{\mathcal{I}_{k}}\|_{2}$ $=\|\sin\Theta(U_{\widehat{\mathcal{I}}_{k}},U_{{\mathcal{I}_{k}}})\|_{2}$
$\leq 2\dim(\mathcal{I}_{k})^{1/2}\left\|P_{\mathcal{J}}\Delta_{k}+\Delta_{\mathcal{J}}P_{\operatorname*{col}(X_{k})}+\Delta_{\mathcal{J}}\Delta_{k}\right\|_{2}$
$=2\dim(\mathcal{I}_{k})^{1/2}\|R_{\mathcal{I}_{k}}\|_{2}.$
A.2. Algorithm implementation
@@
We provide the full implementation of our method in Algorithm 1. Recall from Section 3.3, given the observed data matrices $(Y_{1},Y_{2})$, we first obtain estimates of $\text{col}(\widehat{X}_{1})$ and $\text{col}(\widehat{X}_{2})$ following Gavish & Donoho (2014). Using these estimates, we construct $\widehat{M}=P_{\text{col}(\widehat{X}_{1})}P_{\text{col}(\widehat{X}_{2})}$. We then estimate the rank of the joint subspace $\widehat{r}_{\mathcal{J}}$ by counting the number of singular values in the spectrum of $\widehat{M}$ that exceed $(1-\widehat{\varepsilon}_{1})\vee\lambda_{+}^{1/2}$, where $\widehat{\varepsilon}_{1}$ and $\lambda_{+}$ are defined in equations (4) and (5), respectively. The joint and individual subspaces are estimated according to equation (3).
@@
    Estimate individual subspaces ^$\widehat{\mathcal{I}}_{k}$ as span of the first left $\operatorname*{rank}(\widehat{X}_{k})-\widehat{r}_{\mathcal{J}}$ singular vectors of $P_{\operatorname*{col}(\widehat{X}_{k})}(I-P_{\widehat{\mathcal{J}}})$.
    Output: Joint subspace $\widehat{\mathcal{J}}$ and individual subspaces ^$\widehat{\mathcal{I}}_{1}$ and ^$\widehat{\mathcal{I}}_{2}$; a histogram of the spectrum of $\widehat{M}$ overlaid with $\lambda_{+}^{1/2}$, $1-\widehat{\varepsilon}_{1}$ and the distribution $f(\lambda)$ in (5).
@@
    A.3. Extension to multi-view case
@@
    Suppose our estimate of the joint rank is correct, i.e., $\dim\widehat{\mathcal{J}}=\dim\mathcal{J}$. Define $W_{i_{j}}(\mathcal{S})=P_{\operatorname*{col}(X_{i_{j}})}\mathbb{I}[i_{j}\in\mathcal{S}]+\Delta_{i_{j}}\mathbb{I}[i_{j}\not\in\mathcal{S}]$ and $R_{\mathcal{I}_{k}}:=P_{\mathcal{J}}\Delta_{k}+\Delta_{\mathcal{J}}P_{\operatorname*{col}(X_{k})}+\Delta_{\mathcal{J}}\Delta_{k}$, where $\Delta_{\mathcal{J}}:=P_{\widehat{\mathcal{J}}}-P_{\mathcal{J}}$ and $\Delta_{k}:=P_{{\operatorname*{col}(X_{k})}}-P_{\operatorname*{col}(\widehat{X}_{k})}$ Additionally, let $\mathcal{P}$ be the set of all permutations of $\{1,,2,\dots,K\}$. Then, the deviation of the estimated joint subspace from the true joint subspace is bounded by:
    $\|P_{\widehat{\mathcal{J}}}-P_{\mathcal{J}}\|_{2}\leq\frac{2\dim(\mathcal{J})^{1/2}\left\|\sum_{\mathcal{S}\subset\{1,2,\dots,K\}}\prod_{j=1}^{K}W_{i_{j}}(\mathcal{S})\right\|_{2}}{|\mathcal{P}|\left(1-\frac{1}{|\mathcal{P}|}\sum_{(i_{1},i_{2},\dots,i_{k})\in\mathcal{P}}\left\|P_{\mathcal{I}_{i_{1}}}P_{\mathcal{I}_{i_{2}}}\dots P_{\mathcal{I}_{i_{k}}}\right\|_{2}\right)}.$
@@
    $\|P_{\widehat{\mathcal{I}}_{k}}-P_{\mathcal{I}_{k}}\|_{2}\leq 2\dim(\mathcal{I}_{k})^{1/2}\|R_{\mathcal{I}_{k}}\|_{2}.$
@@
    $\|\sin\Theta(U_{\widehat{\mathcal{J}}},U_{\mathcal{J}})\|_{2}\leq\frac{2\dim(\mathcal{J})^{1/2}\|\widehat{S}-S\|_{2}}{1-\frac{1}{|\mathcal{P}|}\sum_{(i_{1},i_{2},\dots,i_{k})\in\mathcal{P}}\left\|P_{\mathcal{I}_{i_{1}}}P_{\mathcal{I}_{i_{2}}}\dots P_{\mathcal{I}_{i_{k}}}\right\|_{2}}$
@@
    $\|P_{\widehat{\mathcal{J}}}-P_{\mathcal{J}}\|_{2}$ $=\|\sin\Theta(U_{\widehat{\mathcal{J}}},U_{\mathcal{J}})\|_{2}$
    $\leq\frac{2\dim(\mathcal{J})^{1/2}\|\widehat{S}-S\|_{2}}{1-\frac{1}{|\mathcal{P}|}\sum_{(i_{1},i_{2},\dots,i_{k})\in\mathcal{P}}\left\|P_{\mathcal{I}_{i_{1}}}P_{\mathcal{I}_{i_{2}}}\dots P_{\mathcal{I}_{i_{k}}}\right\|_{2}}$
    $\leq\frac{2\dim(\mathcal{J})^{1/2}\left\|\sum_{\mathcal{S}\subset\{1,2,\dots,K\}}\prod_{j=1}^{K}W_{i_{j}}(\mathcal{S})\right\|_{2}}{|\mathcal{P}|\left(1-\frac{1}{|\mathcal{P}|}\sum_{(i_{1},i_{2},\dots,i_{k})\in\mathcal{P}}\left\|P_{\mathcal{I}_{i_{1}}}P_{\mathcal{I}_{i_{2}}}\dots P_{\mathcal{I}_{i_{k}}}\right\|_{2}\right)},$
@@
    A.4. Detailed comparison with related methods

### v2

@@ Appendix A Techinical details
(8) $\left\|P_{{\operatorname*{col}(X_{k})}}-P_{\operatorname*{col}(\widehat{X}_{k})}\right\|_{2}\lesssim\frac{\delta_{k}(n^{1/2}+r_{k}^{1/2})}{\sigma_{r_{k}}(X_{k})}+\frac{\delta_{k}^{2}(n^{1/2}+p_{k}^{1/2})(p_{k}^{1/2}+r_{k}^{1/2})}{\sigma_{r_{k}}(X_{k})^{2}}.$
Rank overestimation case where $\widehat{r}_{k}\geq r_{k}$: As in the previous case, we drop subscript $k$ in the proof for notational convenience. Observe that
$\|P_{{\operatorname*{col}(X)}}\Delta\|_{2}=\|UU^{\top}-UU^{\top}\widehat{U}\widehat{U}^{\top}\|_{2}=\|UU^{\top}\widehat{U}_{\perp}\widehat{U}_{\perp}^{\top}\|=\|U^{\top}\widehat{U}_{\perp}^{\top}\|_{2}.$
Here, $\widehat{U}\in\mathbb{R}^{n\times\widehat{r}}$ is the first $\widehat{r}$ left singular vectors of $Y$ and $\widehat{U}_{\perp}\in\mathbb{R}^{n\times(\mathrm{rank}(Y)-\widehat{r})}$ represents the remaining left singular vectors corresponding to the non-zero singular values. Consider as before the singular value decomposition
$X=[U\ U_{\perp}]\begin{pmatrix}\Sigma_{1}&0\\ 0&0\end{pmatrix}[V\ V_{\perp}]^{\top}.$
Let $Y=X+Z$, and similarly partition $Y$ based on singular value decomposition as follows
$Y=[\widetilde{U}\ \widetilde{U}_{\perp}]\begin{pmatrix}\widetilde{\Sigma}_{1}&0\\ 0&\widetilde{\Sigma}_{2}\end{pmatrix}[\widetilde{V}\ \widetilde{V}_{\perp}$$^{\top}$,
where $\widetilde{U}\in\mathbb{R}^{n\times r}$ and $\widetilde{V}\in\mathbb{R}^{p\times r}$ consists of the first $r$ columns of $\widehat{U}$ and $\widehat{V}$, respectively. Further, $\widetilde{U}_{\perp}$ and $\widetilde{V}_{\perp}$ consist of the remaining columns of $\widehat{U}$ and $\widehat{V}$, respectively. Consider
$\widehat{U}_{\perp}^{\top}U$ $=\widehat{U}_{\perp}^{\top}U\Sigma_{1}V^{\top}V\Sigma_{1}^{-1}=\widehat{U}_{\perp}^{\top}(X)V\Sigma_{1}^{-1}=\widehat{U}_{\perp}^{\top}(Y-Z)V\Sigma_{1}^{-1}$
$=\widehat{U}_{\perp}^{\top}(\widetilde{U}\widetilde{\Sigma}_{1}\widetilde{V}^{\top}+\widetilde{U}_{\perp}\widetilde{\Sigma}_{2}\widetilde{V}_{\perp}^{\top}-Z)V\Sigma_{1}^{-1}$
$=\widehat{U}_{\perp}^{\top}\widetilde{U}_{\perp}\widetilde{\Sigma}_{2}\widehat{V}_{\perp}^{\top}V\Sigma_{1}^{-1}-\widehat{U}_{\perp}^{\top}ZV\Sigma_{1}^{-1}.$
Here, we have used the fact that $\widehat{U}_{\perp}^{\top}\widetilde{U}=0$. Similarly, we can show
$\widehat{V}_{\perp}^{\top}V=\widehat{V}_{\perp}^{\top}\widetilde{V}\widetilde{\Sigma}_{2}\widehat{U}_{\perp}^{\top}U\Sigma_{1}^{-1}-\widehat{V}_{\perp}^{\top}Z^{\top}U\Sigma_{1}^{-1}.$
Combining the above two displays gives
$\widehat{U}_{\perp}^{\top}U$ $=\widehat{U}_{\perp}^{\top}\widetilde{U}_{\perp}\widetilde{\Sigma}_{2}\widehat{V}_{\perp}^{\top}V\widetilde{\Sigma}_{2}\widehat{U}_{\perp}^{\top}U\Sigma_{1}^{-2}-\widehat{U}_{\perp}^{\top}\widetilde{U}_{\perp}\widetilde{\Sigma}_{2}\widehat{V}_{\perp}^{\top}Z^{\top}U\Sigma_{1}^{-2}-\widehat{U}_{\perp}^{\top}ZV\Sigma_{1}^{-1}.$
Therefore, using triangle inequality, the sub-multipicative property of spectral norms, that $\|\widehat{U}_{\perp}^{\top}\widetilde{U}_{\perp}\|_{2}\leq 1$, and that $\|\Sigma_{1}^{-2}\|_{2}=\frac{1}{\sigma_{r}(X)^{2}}$, we obtain:
$\|\widehat{U}_{\perp}^{\top}U\|_{2}\leq\|\widehat{U}_{\perp}^{\top}U\|_{2}\|\widetilde{\Sigma}_{2}\|_{2}^{2}\frac{1}{\sigma_{r}(X)^{2}}+\|\widetilde{\Sigma}_{2}\|_{2}\frac{1}{\sigma_{r}(X)^{2}}\|\widehat{V}_{\perp}^{\top}Z^{\top}U\|_{2}+\frac{1}{\sigma_{r}(X)}\|\widehat{U}_{\perp}^{\top}ZV\|_{2}.$
As in the previous case, appealing to Weyl’s inequality and using that $X$ is rank-$r$, we have that $\|\widehat{\Sigma}_{2}\|_{2}\leq\|Z\|_{2}$. Following the same logic as the correct rank case, we conclude that:
$\|P_{{\operatorname*{col}(X_{k})}}\Delta_{k}\|_{2}=\|U^{\top}\widehat{U}_{\perp}^{\top}\|_{2}\lesssim\frac{\delta_{k}(n^{1/2}+r_{k}^{1/2})}{\sigma_{r_{k}}(X_{k})}+\frac{\delta_{k}^{2}(n^{1/2}+p_{k}^{1/2})(p_{k}^{1/2}+r_{k}^{1/2})}{\sigma_{r_{k}}(X_{k})^{2}}.$
 of Theorem 3.3.
We consider the bounding of $\varepsilon_{1}$, $\varepsilon_{2}$, and the estimation accuracy of subspaces separately.
Bounding $\varepsilon_{1}$: Recall,
$\begin{aligned} \varepsilon_{1}&:=\left\|P_{{\operatorname*{col}(X_{1})}}\left(\Delta_{1}+\Delta_{2}+\Delta_{1}\Delta_{2}\right)P_{{\operatorname*{col}(X_{2})}}\right\|_{2},\end{aligned}$
with $\Delta_{k}:=P_{{\operatorname*{col}(X_{k})}}-P_{\operatorname*{col}(\widehat{X}_{k})}$. By triangle inequality, symmetry of projection matrices, and sub-multipicative property of the spectral norm, we have that:
$\varepsilon_{1}\leq\left\|P_{{\operatorname*{col}(X_{1})}}\Delta_{1}\right\|_{2}+\left\|P_{{\operatorname*{col}(X_{2})}}\Delta_{2}\right\|+\left\|P_{{\operatorname*{col}(X_{1})}}\Delta_{1}\right\|_{2}\left\|P_{{\operatorname*{col}(X_{2})}}\Delta_{2}\right\|_{2}.$
Under the assumption $\widehat{r}_{k}\geq r_{k}$, we appeal to Lemma A.5 to conclude that with probability greater than $1-\mathcal{O}(\exp\{-\max(p_{\min},n)\})$:
$\varepsilon_{1}\lesssim\frac{\delta_{\max}(n^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)}+\frac{\delta_{\max}^{2}(n^{1/2}+p_{\max}^{1/2})(p_{\max}^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)^{2}}.$
Here, we have used that $\|\mathcal{P}_{\operatorname*{col}(X_{1})}\Delta_{1}\|_{2}\|\mathcal{P}_{\operatorname*{col}(X_{2})}\Delta_{2}\|_{2}\lesssim\max_{k}\|\mathcal{P}_{\operatorname*{col}(X_{k})}\Delta_{k}\|_{2}$ for $k=1,2$ due to Assumption 2.
Bounding $\varepsilon_{2}$: Recall,
$\begin{aligned} \varepsilon_{2}&:=\left\|P_{{\operatorname*{col}(X_{1})}}\Delta_{2}+\Delta_{1}P_{{\operatorname*{col}(X_{2})}}+\Delta_{1}\Delta_{2}\right\|_{2}.\end{aligned}$
Using triangle inequality and the sub-multipicative property of the spectral norm, we have that:
$\varepsilon_{2}\leq\|\Delta_{1}\|_{2}+\|\Delta_{2}\|_{2}+\|\Delta_{1}\|_{2}\|\Delta_{2}\|_{2}.$
Under the correctly specified rank $\widehat{r}_{k}=r_{k}$, we appeal to Lemma A.5 to conclude that with probability greater than $1-\mathcal{O}(\exp\{-\max(p_{\min},n)\})$:
$\varepsilon_{2}\lesssim\frac{\delta_{\max}(n^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)}+\frac{\delta_{\max}^{2}(n^{1/2}+p_{\max}^{1/2})(p_{\max}^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)^{2}}.$
Here, we have used the property that $\|\Delta_{k}\|_{2}^{2}\lesssim\|\Delta_{k}\|_{2}$ for $k=1,2$ due to Assumption 2.
Bounding subspace estimation errors: Since $\widehat{r}_{J}=r_{J}$ and $\widehat{r}_{k}=r_{k}$, appealing to Theorem 3.2, we have that:
$\begin{aligned} \left\|P_{\mathcal{J}}-P_{\widehat{\mathcal{J}}}\right\|_{2}&\leq\frac{\left\|R_{\mathcal{J}}+R^{\top}_{\mathcal{J}}\right\|_{2}}{1-\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}},\\ \left\|P_{\mathcal{I}_{k}}-P_{\widehat{\mathcal{I}}_{k}}\right\|_{2}&\leq{2}\|R_{\mathcal{I}_{k}}\|_{2},\end{aligned}$
where ${R}_{\mathcal{J}}:=P_{\operatorname*{col}(X_{1})}\Delta_{2}+\Delta_{1}P_{\operatorname*{col}(X_{2})}+\Delta_{1}\Delta_{2}$ and $R_{\mathcal{I}_{k}}:=P_{\mathcal{J}^{\perp}}\Delta_{k}-\Delta_{\mathcal{J}}P_{\operatorname*{col}(X_{k})}+\Delta_{\mathcal{J}}\Delta_{k}$ and $\Delta_{\mathcal{J}}:=P_{\mathcal{J}}-P_{\widehat{\mathcal{J}}}$. Using triangle inequality, sub-multipicative property of spectral norm, and that the spectral norm of a projection matrix is one, we have that:
$\begin{aligned} \|{R}_{\mathcal{J}}\|_{2}&\leq\|\Delta\|_{2}+\|\Delta\|_{1}+\|\Delta_{1}\|_{2}\|\Delta_{2}\|_{2},\\ \|{R}_{\mathcal{I}_{k}}\|_{2}&\leq\|\Delta_{k}\|_{2}+\|\Delta_{\mathcal{J}}\|_{2}+\|\Delta_{k}\|_{2}\|\Delta_{\mathcal{J}}\|_{2}.\end{aligned}$
Appealing to Lemma A.5, we have with probability greater than $1-\mathcal{O}(\exp\{-\max(p_{\min},n)\})$:
@@
$\begin{aligned} \left\|P_{\mathcal{J}}-P_{\widehat{\mathcal{J}}}\right\|_{2}\lesssim\frac{1}{1-\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}}\left\{\frac{\delta_{\max}(n^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)}+\frac{\delta_{\max}^{2}(n^{1/2}+p_{\max}^{1/2})(p_{\max}^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)^{2}}\right\},\\ \left\|P_{\mathcal{I}_{k}}-P_{\widehat{\mathcal{I}_{k}}}\right\|_{2}\lesssim\frac{1}{1-\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}}\left\{\frac{\delta_{\max}(n^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)}+\frac{\delta_{\max}^{2}(n^{1/2}+p_{\max}^{1/2})(p_{\max}^{1/2}+r_{\max}^{1/2})}{\sigma_{r}(X)^{2}}\right\}.\end{aligned}$
A.4. Algorithm implementation
@@
We provide the full implementation of our method in Algorithm 1. Recall from Section 4, given the observed data matrices $(Y_{1},Y_{2})$, we first obtain estimates of $\text{col}(\widehat{X}_{1})$ and $\text{col}(\widehat{X}_{2})$ following Gavish & Donoho (2014). Using these estimates, we construct $\widehat{M}=P_{\text{col}(\widehat{X}_{1})}P_{\text{col}(\widehat{X}_{2})}$. We then estimate the rank of the joint subspace $\widehat{r}_{\mathcal{J}}$ by counting the number of singular values in the spectrum of $\widehat{M}$ that exceed $(1-\widehat{\varepsilon}_{1})\vee\lambda_{+}^{1/2}$, where $\widehat{\varepsilon}_{1}$ and $\lambda_{+}$ are defined in equations (4) and (5), respectively. The joint and individual subspaces are estimated according to equation (3).
@@
[…]
    Output: Joint subspace $\widehat{\mathcal{J}}$ and individual subspaces $\widehat{\mathcal{I}}_{1}$ and $\widehat{\mathcal{I}}_{2}$; a histogram of the spectrum of $\widehat{M}$ overlaid with $\lambda_{+}^{1/2}$, $1-\widehat{\varepsilon}_{1}$ and the distribution $f(\lambda)$ in (5).
@@
    A.5. Extension to multi-view case
@@
    Suppose our estimate of the joint rank is correct, i.e., $\dim\widehat{\mathcal{J}}=\dim\mathcal{J}$. Define $W_{i_{j}}(\mathcal{S})=P_{\operatorname*{col}(X_{i_{j}})}\mathbb{I}[i_{j}\in\mathcal{S}]+\Delta_{i_{j}}\mathbb{I}[i_{j}\not\in\mathcal{S}]$ and $R_{\mathcal{I}_{k}}:=P_{\mathcal{J}}\Delta_{k}-\Delta_{\mathcal{J}}P_{\operatorname*{col}(X_{k})}+\Delta_{\mathcal{J}}\Delta_{k}$, where $\Delta_{\mathcal{J}}:=P_{\mathcal{J}}-P_{\widehat{\mathcal{J}}}$ and $\Delta_{k}:=P_{{\operatorname*{col}(X_{k})}}-P_{\operatorname*{col}(\widehat{X}_{k})}$ . Additionally, let $\mathcal{P}$ be the set of all permutations of $\{1,,2,\dots,K\}$. Then, the deviation of the estimated joint subspace from the true joint subspace is bounded by:
    $\|P_{\mathcal{J}}-P_{\widehat{\mathcal{J}}}\|_{2}\leq\frac{{2}\left\|\sum_{\mathcal{S}\subset\{1,2,\dots,K\}}\prod_{j=1}^{K}W_{i_{j}}(\mathcal{S})\right\|_{2}}{|\mathcal{P}|\left(1-\frac{1}{|\mathcal{P}|}\sum_{(i_{1},i_{2},\dots,i_{k})\in\mathcal{P}}\left\|P_{\mathcal{I}_{i_{1}}}P_{\mathcal{I}_{i_{2}}}\dots P_{\mathcal{I}_{i_{k}}}\right\|_{2}\right)}.$
@@
    $\|P_{\mathcal{I}_{k}}-P_{\widehat{\mathcal{I}}_{k}}\|_{2}\leq{2}\|R_{\mathcal{I}_{k}}\|_{2}.$
@@
    $\|\sin\Theta(U_{\mathcal{J}},U_{\widehat{\mathcal{J}}})\|_{2}\leq\frac{{2}\|S-\widehat{S}\|_{2}}{1-\frac{1}{|\mathcal{P}|}\sum_{(i_{1},i_{2},\dots,i_{k})\in\mathcal{P}}\left\|P_{\mathcal{I}_{i_{1}}}P_{\mathcal{I}_{i_{2}}}\dots P_{\mathcal{I}_{i_{k}}}\right\|_{2}}$
@@
    $\|P_{\mathcal{J}}-P_{\widehat{\mathcal{J}}}\|_{2}$ $=\|\sin\Theta(U_{\mathcal{J}},U_{\widehat{\mathcal{J}}})\|_{2}$
    $\leq\frac{{2}\|S-\widehat{S}\|_{2}}{1-\frac{1}{|\mathcal{P}|}\sum_{(i_{1},i_{2},\dots,i_{k})\in\mathcal{P}}\left\|P_{\mathcal{I}_{i_{1}}}P_{\mathcal{I}_{i_{2}}}\dots P_{\mathcal{I}_{i_{k}}}\right\|_{2}}$
    $\leq\frac{{2}\left\|\sum_{\mathcal{S}\subset\{1,2,\dots,K\}}\prod_{j=1}^{K}W_{i_{j}}(\mathcal{S})\right\|_{2}}{|\mathcal{P}|\left(1-\frac{1}{|\mathcal{P}|}\sum_{(i_{1},i_{2},\dots,i_{k})\in\mathcal{P}}\left\|P_{\mathcal{I}_{i_{1}}}P_{\mathcal{I}_{i_{2}}}\dots P_{\mathcal{I}_{i_{k}}}\right\|_{2}\right)},$
@@
    A.6. Detailed comparison with related methods